OpenStudy (anonymous):

The sum of a number and its square is 110. Find the number. I'll give you a medal :D

OpenStudy (anonymous):

10

OpenStudy (anonymous):

I know the answer but I need the steps... ;)

OpenStudy (anonymous):

Simple x^2 = 100 + 10 = 110. Do you want the algebra?

OpenStudy (anonymous):

Yes, please :)

OpenStudy (tkhunny):

Seriously? We're just guessing and lucky because this one is ralatively obviouss. Rule #1 - Name Things! N - The Number We Seek The sum of a number and its square \(N + N^{2}\) You need to see that translation is 110 \(N + N^{2} = 110\) Solve the quadratic equation. There are TWO solutions.

OpenStudy (anonymous):

It's also -11

OpenStudy (anonymous):

Thank you very much :D

OpenStudy (anonymous):

(x-10)(x+11) = 0

OpenStudy (anonymous):

Then x -10 is 10 x + 11 is -11